OpenStudy (anonymous):

A trophy 8 feet tall weighs 10 ounces. How much would a similar trophy 6 feet tall weigh? Round your answer to the nearest tenth.

OpenStudy (p0sitr0n):

8/6 =k V1/V2 = k^3 V2 =4.22

OpenStudy (anonymous):